The following reaction has an activation energy of 262 kJ/mol.

C4H8(g)-------> 2C2H4(g)

At 600.0 K the rate constant is 6.1
Answer #1

ln(k2/k1)=Ea/R( 1/T1 - 1/T2)
Ea/R(1-T1 - 1/T2) = 26200 J*Mol^-1 / 8.314 J*Mol*K^-1 _1/600K - 1/865K) = 16.09
ln (k2/k1) = 16.09
k2/k1 = e^16.09 = 9.72 *10 ^6
k2 =9.72*10^6k1 = (9.72*10^6) (6.1*10^-8 s^-1 = 0.59s^-1
Answer #2

  Given:
T1 = 600 K
T2 = 875 K
R = 8.314 J/mol K
K1 = 6.1 x 10^-8
K2 = ??
Ea = 262 kj/mol which you convert to J/mol because R is in J/mol so...
Ea = 262000

Now put in the values using the Arrhenius equation:
ln(k2/k1) = (Ea / R) x (1/T1